# What should a neuron aim for? Designing local objective functions based on information theory

### v4

@@ A.1 Pseudocode for Training Algorithm (Setup 1)
Note that, as originally proposed by Kay (1994) and utilized by Makkeh et al. (2025), gradients for the incoming weights are only computed with respect to the conditional probability $p(y|f,c,l)$ (or $p(y|f,c)$) of the output $Y$, since the empirical probability mass functions $p(f,c,l)$ in Algorithm 2 and $p(f,c)$ in Algorithm 3 are not differentiable. While we currently have no strong a priori argument why the empirical probability mass functions can be assumed constant, the learning of PID over training and the good task performance when using these gradients in experiments provides an ex post justification for this procedure.
@@ A.3 Compute Resources
For the figures in this paper, we performed a total of seven hyperparameter optimizations: Three TPE optimizations and three CMA-ES optimizations of MNIST objective function parameter and a single CMA-ES optimization for the CIFAR-10 task. Each of these optimizations took $\leq 12$ hours to compute on a HPC cluster node with $32$ cpu cores and two NVIDIA A100 GPUs with 404040\,40GB of VRAM each.
@@ A.4 Implementation Details and Model Parameters
The parameters used to train the models shown in Fig. 3 are listed in Table 1. The backpropagation model was trained to optimize the cross-entropy loss between the prediction and the true label.
Table 1: Model and training parameters for the models shown in Fig. 3. *The effect of different batch sizes is analyzed in Fig. 10. **The effect of the number of bins is analyzed in Fig. 11
@@ A.5 Model Dynamics During Training
where $\boldsymbol{w}_{F}$ corresponds to feedforward weights of a single neuron of the hidden layer. The median value of $D_{c}^{(t)}$ for the hidden neurons of a trivariate model are shown in Fig. 5. While the self-cosine distance consistently increases with layer size in the dense setups, it does not increase in the sparse setup after a layer size of 100 neurons, which is also the number of lateral connections in all of the larger sparse layers.
@@ A.6 Optimized Objective Functions for the Trivariate Neurons
To support the trivariate objective function that was obtained using hyperparameter optimization, we repeated the optimizations three times for both, the TPE sampler and the CMA-ES sampler, with parameter values being limited to the interval $\gamma_{i}\in[-1,1]$. The resulting objective functions are illustrated in Fig. 6. While the objective functions that were optimized with the TPE sampler included all $\gamma$ parameters, the optimization with the CMA-ES sampler was only performed for the parameters that describe PID-atoms while the parameter for the residual entropy was set to 0. This was done because we observed a strong correlation between the parameter for the unique information of the feedforward input and the residual entropy which both had to be small. In the figure, the respective values are set 0. In the figure, we also illustrate the mean decrease impurity score which is a measure for the importance of a parameter.
@@ A.7 Trivariate Consistency Equations
To improve the clarity and intuition for the atoms of the partial information decomposition with three source variables, Fig. 7 shows a larger version of decomposition diagram already shown in Fig. 2.C. As a guide for the reader, the meaning of the different information atoms are listed in Table 2. Additionally, Table 3 explicitly lists the goal parameter values found via optimization together with their interpretation.
Figure 7: Larger version of decomposition diagram already shown in Fig. 2.
@@
$\gamma_{\{F\}\{L\}}$ redundancy between $F$ and $L$0.99 minimize information that is redundant with other neurons, but NOT provided by the context—and thus not relevant for the task
@@
$\gamma_{\{FC\}\{FL\}}$ redundancy between synergies requiring $F$0.97 minimize synergistic information that requires $F$ and either $C$ or $L$ to be recovered
Figure 8: An overview over the training and test performances of the heuristic (left) and optimized (right) models with different model structures, neuron models and objective functions as well as the performances of some baseline models. In the bivariate models, only one of the non driving inputs (context or lateral) and the neurons optimized the heuristic objective functions illustrated in Fig. 2.D. The corresponding optimized objectives were optimized with the CMA-ES sampler similar to the trivariate objective function.
[…]
@@ A.9 Recurrence of the Hidden Layer
To perform an information decomposition which includes the information of the lateral activity, we include lateral connections within the hidden layer as illustrated in Fig. 3.A. As described in Section A.1, we only perform two iterations within the hidden layer i.e.,one iteration to compute the activity of the neurons within the layer and a second iteration to include this activity in the final output of the neurons. Therefore, the lateral activity contributes to the neurons output. Due to the stochasticity of the neurons, the lateral input will never fully converge. The continuous output probabilities converge already after the second single iteration with only minor changes in the output probabilities of the order of $10^{-4}$. This is due to the small influence of the lateral input on the output which is limited to $0.1\,\sigma(2FL)$ i.e., a maximum change of ±$\pm 0.1$ in the activation.

### v5

@@ A.1 Pseudocode for Training Algorithm (Setup 1)
Note that, as originally proposed by Kay (1994) and utilized by Makkeh et al. (2025), gradients for the incoming weights are only computed with respect to the conditional probability $p(y|f,c,l)$ (or $p(y|f,c)$) of the output $Y$, since the empirical probability mass functions $p(f,c,l)$ in Algorithm˜2 and $p(f,c)$ in Algorithm˜3 are not differentiable. While we currently have no strong a priori argument why the empirical probability mass functions can be assumed constant, the learning of PID over training and the good task performance when using these gradients in experiments provides an ex post justification for this procedure.
@@ A.3 Compute Resources
For the figures in this paper, we performed a total of seven hyperparameter optimizations: Three TPE optimizations and three CMA-ES optimizations of MNIST objective function parameter and a single CMA-ES optimization for the CIFAR-10 task. Each of these optimizations took $\leq 12$ hours to compute on a HPC cluster node with $32$ cpu cores and two NVIDIA A100 GPUs with 4040\,GB of VRAM each.
@@ A.4 Implementation Details and Model Parameters
The parameters used to train the models shown in Fig.˜3 are listed in Table˜1. The backpropagation model was trained to optimize the cross-entropy loss between the prediction and the true label.
Table 1: Model and training parameters for the models shown in Fig.˜3. *The effect of different batch sizes is analyzed in Fig.˜10. **The effect of the number of bins is analyzed in Fig.˜11
@@ A.5 Model Dynamics During Training
where $\boldsymbol{w}_{F}$ corresponds to feedforward weights of a single neuron of the hidden layer. The median value of $D_{c}^{(t)}$ for the hidden neurons of a trivariate model are shown in Fig.˜5. While the self-cosine distance consistently increases with layer size in the dense setups, it does not increase in the sparse setup after a layer size of 100 neurons, which is also the number of lateral connections in all of the larger sparse layers.
@@ A.6 Optimized Objective Functions for the Trivariate Neurons
To support the trivariate objective function that was obtained using hyperparameter optimization, we repeated the optimizations three times for both, the TPE sampler and the CMA-ES sampler, with parameter values being limited to the interval $\gamma_{i}\in[-1,1]$. The resulting objective functions are illustrated in Fig.˜6. While the objective functions that were optimized with the TPE sampler included all $\gamma$ parameters, the optimization with the CMA-ES sampler was only performed for the parameters that describe PID-atoms while the parameter for the residual entropy was set to 0. This was done because we observed a strong correlation between the parameter for the unique information of the feedforward input and the residual entropy which both had to be small. In the figure, the respective values are set 0. In the figure, we also illustrate the mean decrease impurity score which is a measure for the importance of a parameter.
@@ A.7 Trivariate Consistency Equations
To improve the clarity and intuition for the atoms of the partial information decomposition with three source variables, Fig.˜7 shows a larger version of decomposition diagram already shown in Fig.˜2.C. As a guide for the reader, the meaning of the different information atoms are listed in Table˜2. Additionally, Table˜3 explicitly lists the goal parameter values found via optimization together with their interpretation.
Figure 7: Larger version of decomposition diagram already shown in Fig.˜2.
@@
$\gamma_{\{F\}\{L\}}$ redundancy between $F$ and $L$ -0.99 minimize information that is redundant with other neurons, but NOT provided by the context—and thus not relevant for the task
@@
$\gamma_{\{FC\}\{FL\}}$ redundancy between synergies requiring $F$ -0.97 minimize synergistic information that requires $F$ and either $C$ or $L$ to be recovered
Figure 8: An overview over the training and test performances of the heuristic (left) and optimized (right) models with different model structures, neuron models and objective functions as well as the performances of some baseline models. In the bivariate models, only one of the non driving inputs (context or lateral) and the neurons optimized the heuristic objective functions illustrated in Fig.˜2.D. The corresponding optimized objectives were optimized with the CMA-ES sampler similar to the trivariate objective function.
Figure 9: To estimate the importance of the different parameters of the trivariate objective function, we modified individual parameters of the trivariate objective function. The performance change induced by setting the parameters to 0 was already shown in Fig.˜4.B and was used for the results shown in Fig.˜4.C. Additionally, we changed the individual goal parameter relative to the parameter strength ($\pm 10\%)$, but also in an absolute manner ($\pm 0.1$) to get a better intuition about the importance of the fine tuning of the parameters.
@@ A.9 Recurrence of the Hidden Layer
To perform an information decomposition which includes the information of the lateral activity, we include lateral connections within the hidden layer as illustrated in Fig.˜3.A. As described in Section˜A.1, we only perform two iterations within the hidden layer i.e.,one iteration to compute the activity of the neurons within the layer and a second iteration to include this activity in the final output of the neurons. Therefore, the lateral activity contributes to the neurons output. Due to the stochasticity of the neurons, the lateral input will never fully converge. The continuous output probabilities converge already after the second single iteration with only minor changes in the output probabilities of the order of $10^{-4}$. This is due to the small influence of the lateral input on the output which is limited to $0.1\,\sigma(2FL)$ i.e., a maximum change of $\pm 0.1$ in the activation.